Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours. In the most recent month, 120,000 items were shipped to customers using 2,300 direct labor-hours. The company incurred a total of $7,360 in variable overhead costs.

According to the company’s standards, 0.02 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $3.25 per direct labor-hour.

Required:

1. What is the standard labor-hours allowed (SH) to ship 120,000 items to customers?

2. What is the standard variable overhead cost allowed (SH × SR) to ship 120,000 items to customers?

3. What is the variable overhead spending variance?

4. What is the variable overhead rate variance and the variable overhead efficiency variance?
Answer #1

1. Standard Labor hours allowed = 120000 x 0.02 = 2400 hours

2. Standard Variable Overhead Cost = 2400 x $3.25 = $7800

3. Variable Overhead Spending Variance = Actual Overhead - Standard Overhead
= $7360 - $7800 = $440 (F)

4.
Variable Overhead Rate Variance = Actual Overhead - Standard Rate x Actual hours
= $7360 - (2300 x $3.25) = $115 (F)

Variable Overhead Efficiency Variance = (Actual hours - Standard hours) x Standard Rate
= (2300 - 2400) x $3.25 = $325 (F)